# Improving Coverage in Combined Prediction Sets with Weighted p-values

### v1

@@ 5 Extending to data-dependent weights
The prediction set $\widehat{C}_{\alpha}^{\text{data-avg}}(X_{n+1})$ defined by the weighted average $p_{\text{all}}$ (7) has the following coverage guarantee:
$\mathbb{P}\left\{Y_{n+1}\in\widehat{C}^{\text{data-avg}}_{\alpha}(X_{n+1})\right\}\geq 1-\left(\alpha\mathbb{E}[\widehat{m}^{*}]+\varepsilon+\delta\right)\quad\forall\alpha\in(0,1),$ (12)
where $\widehat{m}^{*}$ is defined in (11).
In §B.1.1, we explore how the size of $S_{\text{merge}}$ affects $\widehat{F}^{\text{cons}}_{P_{\text{all}}}$ and therefore, the quality of the correction $\widehat{m}^{*}$ and the tightness of the coverage guarantee. We find that in practice, a modest number of samples ($<200$) is sufficient to achieve accurate coverage. We summarize our procedure in Algorithm 1.
@@
Data-dependent weights allow the influence of each model to be adjusted based on how well it performs for a specific data point. This approach enables the construction of aggregated prediction sets that are tailored to the characteristics of the given data, which can be viewed as a form of locally conditional coverage. To be clear, true X-conditional coverage—where the coverage guarantee is conditioned on the current input—is impossible without additional distributional assumptions [33, 34]. However, we demonstrate in §6 that data-dependent weights allow us to create aggregated prediction sets with data-adaptive coverage that can greatly improve conditional validity in practice.
@@
Note that (13) is not without precedent: for the $1-2\alpha$ coverage guarantees of aggregation methods like cross-conformal and jackknife+, $\alpha^{\prime}=0.5$ is the highest significance level of interest.
@@ 6 Application: Mixture-of-Experts
Refer to captionRefer to caption
Refer to captionRefer to caption
@@
The routing network’s ability to learn data-dependent weights for each expert makes the MoE framework a natural setting for applying our method of aggregating prediction sets by weighted p-values. We collect the routing weights into the simplex-valued vector $W(x):=(W_{1}(x),\dots,W_{K}(x))$; then, for a new input $X_{n+1}$, the routing network outputs the weight vector $\mathbf{W}^{(n+1)}:=W(X_{n+1})$, whose $k$th component we write as $W^{(n+1)}_{k}$. Denote the p-value function of the $k$th expert as $\widehat{p}_{k}$. The weighted aggregate p-value function for some label $y$ is then
@@
At a significance level $\alpha$, the associated prediction set is
$\widehat{C}_{\alpha}^{\text{data-avg}}(X_{n+1})=\{y\in\mathcal{Y}:p_{\text{all}}(X_{n+1},y;\mathbf{W}^{(n+1)})>\alpha\}.$
Intuitively, the weights assigned by the routing network adjust the p-value contribution of each expert by how predictive they are for a given point. Data-dependent weights enable data-adaptive coverage, which is more locally robust than split conformal with the full MoE (15) as a predictor (Figure 1).
@@
We compare against conformal quantile regression (CQR) [35] for experiments evaluating the local validity of our method. Weighted aggregation can be applied on top of conformal quantile regression, so we also evaluate a hybrid method combining the two. To assess performance, we consider marginal coverage, worst-slice (WS) coverage [36], and prediction set size.
@@
Results
In Figure 2, we compare split conformal to weighted aggregation using both absolute residual scores and CQR scores. From the marginal coverage plots (left column), we see that split conformal consistently achieves coverage closest to the nominal $1-\alpha$ level, while the weighted aggregation methods tend to overcover. However, the WS coverage plots (middle column) highlight a well-known limitation of split conformal: it does not maintain coverage uniformly across the data, significantly undercovering on the WS slab. While CQR scores improve split conformal WS coverage on some datasets (Airfoil and Communities), they are still insufficient to close the coverage gap.
Weighted aggregation significantly improves WS coverage, coming close to or meeting the nominal level on all datasets. In addition, the gap between marginal and WS coverage is notably smaller for weighted aggregation than for split conformal, indicating that weighted aggregation provides more uniform coverage, including in challenging regions of the data. These results suggest that weighted aggregation is useful in applications where local validity is a priority.
@@ 7 Conclusion
The asymmetric (weighted) aggregation of prediction sets is a flexible generalization of the more standard symmetric aggregation. Intuitively, coverage for this case should not be constant, but should depend on the distribution of weights. The results of Vovk and Wang [18] enable us to achieve this for conformal prediction, so that for prior weights using expert information, we improve the coverage guarantee when the distribution of weights is sufficiently asymmetric. In addition, we propose a method for data-dependent weights (e.g. weights learned from data) where coverage depends on the weight distribution. We find that allowing for data-dependent weights provides a form of conditional coverage, which we verify in experiments on WS coverage and coverage across demographic groups.
@@ Acknowledgements
GW and RC are partially supported by a ONR MURI grant N00014-20-1-2781. DP is partially supported by an Amazon AI2AI fellowship. DP, SS, and AL are partially supported by the Gordon and Betty Moore Foundation grant #12128. AL is also supported by an Amazon Research Award.
@@ Appendix A Proofs and additional theoretical details
A.2 Proof of Proposition 5.1
Proposition 2 follows from the definition of the correction factor $m^{*}$ (8).
We define $m^{*}$ to be the smallest positive value such that $m^{*}P_{\text{all}}$ is a p-variable, meaning that
$\mathbb{P}\left\{m^{*}P_{\text{all}}\leq\alpha\right\}\leq\alpha.$
We can rearrange this condition to be
@@
$\mathbb{P}\left\{P_{\text{all}}\leq\alpha^{\prime}\right\}\leq m^{*}\alpha^{\prime},$
or, equivalently,
@@
$\widehat{C}^{\text{data-avg}}_{\alpha}(X_{n+1})=\left\{y\in\mathcal{Y}:p_{\text{all}}(X_{n+1},y;\mathbf{W}^{(n+1)})>\alpha\right\}$
satisfies the coverage guarantee
$\mathbb{P}\left\{Y_{n+1}\in\widehat{C}^{\text{data-avg}}_{\alpha}(X_{n+1})\right\}\geq 1-m^{*}\alpha.$
Note that other transformations of $P_{\text{all}}$ can also yield valid p-variables while preserving the relative weighting. For example, if $P_{\text{all}}$ is shifted rather than scaled, then the condition
@@
leads to a coverage guarantee of $1-(m^{\prime}+\alpha)$. We focus on the scale correction factor $m^{*}$ because it aligns with the framework of Vovk and Wang [18] and works well in practice.
A.3 Proof of Proposition 5.2
@@
$\mathbb{P}\left\{\widehat{m}^{*}P_{\text{all}}\leq\alpha\mid S_{\text{merge}}\right\}\leq\alpha+\varepsilon,$
or
$\mathbb{P}\left\{P_{\text{all}}\leq\alpha^{\prime}\mid S_{\text{merge}}\right\}\leq\widehat{m}^{*}\alpha^{\prime}+\varepsilon$
by a change of variable.
Recall that $P_{\text{all}}:=p_{\text{all}}(X,Y,\mathbf{W})$. Under exchangeability, the prediction set
$\widehat{C}^{\text{data-avg}}_{\alpha}(X_{n+1})=\left\{y\in\mathcal{Y}:p_{\text{all}}(X_{n+1},y;\mathbf{W}^{(n+1)})>\alpha\right\}$
@@
$\mathbb{P}\left\{Y_{n+1}\notin\widehat{C}^{\text{data-avg}}_{\alpha}(X_{n+1})\mid S_{\text{merge}}\right\}\leq\widehat{m}^{*}\alpha+\varepsilon.$
@@
$\text{MC}=\mathbbm{1}\left\{Y_{n+1}\notin\widehat{C}^{\text{data-avg}}_{\alpha}(X_{n+1})\right\}.$
@@
This gives us the coverage guarantee
$\mathbb{P}\left\{Y_{n+1}\in\widehat{C}^{\text{data-avg}}_{\alpha}(X_{n+1})\right\}\geq 1-\left(\alpha\mathbb{E}[\widehat{m}^{*}]+\varepsilon+\delta\right).$
Note that the same argument applies directly if we replace the empirical CDF $\widehat{F}_{P_{\text{all}}}$ by the conservative version $\widehat{F}_{P_{\text{all}}}^{\text{cons}}$ (10) because the two differ by at most $(|\mathcal{S}_{\text{merge}}|+1)^{-1}$, a deterministic offset that can simply be absorbed into the DKW tolerance by replacing $\varepsilon$ with $\varepsilon+(|\mathcal{S}_{\text{merge}}|+1)^{-1}$.
@@ A.5 Computing $\widehat{m}^{*}$ from an empirical CDF
$\lambda\in\{m>0:\mathbb{P}\{mP_{\text{all}}\leq\alpha\}\leq\alpha\leavevmode\nobreak\ \forall\alpha\in(0,1)\}=\{m>0:F_{P_{\text{all}}}\left(\alpha/m\right)\leq\alpha\leavevmode\nobreak\ \forall\alpha\in(0,1)\}.$
@@
$m^{*}=\inf\{m>0:F_{P_{\text{all}}}\left(\alpha/m\right)\leq\alpha\leavevmode\nobreak\ \forall\alpha\in(0,1)\}=\sup_{\delta>0}{\frac{F_{P_{\text{all}}}(\delta)}{\delta}}.$ (19)
@@
$\widehat{F}_{P_{\text{all}}}(\alpha)=\frac{\sum_{i\in S_{\text{merge}}}\mathbbm{1}\left\{p_{\text{all}}(X_{i},Y_{i};\mathbf{W}^{(i)})\leq\alpha\right\}}{|S_{\text{merge}}|},$
@@ B.1 Synthetic data
Figure 3 illustrates how the size of the merging set $|S_{\text{merge}}|$ impacts mean coverage. On the left plot, we use the $m^{\dagger}$ correction (13) for WA targeted (that is, $(0,\alpha)$ aggregation), and we compare different ways of assigning features to experts. On the right plot, we use a non-overlapping feature assignment, and we compare the different variants of weighted aggregation. Our results show that larger $|S_{\text{merge}}|$ leads to tighter coverage, with 160 samples being sufficient to achieve $<3\%$ overcoverage for most feature assignment methods. Interestingly, configurations where experts have no overlapping features tend to overcover the most. We explore the effect of feature assignment in the next section.
@@ B.2.2 Local validity over demographic groups
Figures 7 and 8 compare split conformal with weighted aggregation variants in terms of coverage and prediction set size across demographic groups. Unlike Figures 2 and 6, where local validity is assessed via WS coverage, these experiments evaluate local validity in terms of consistency across demographic groups. Thus, we display group-specific performance for each method, which we group further by the two types of nonconformity score, absolute residual and CQR score.
The demographic groups in Figures 7 and 8 represent communities where a particular racial demographic is in the top $p$-percentile of representation [38]. Figure 7 shows results for $p=50$, and Figure 8 for $p=70$.

### v2

@@ 5 Extending to data-dependent weights
The prediction set from thresholding $p_{\text{all}}$ is $\widehat{C}_{\alpha}^{\text{unscaled}}(X_{n+1})=\{y:p_{\text{all}}(X_{n+1},y)>\alpha\}$, and it satisfies the coverage guarantee
$\mathbb{P}\{Y_{n+1}\in\widehat{C}^{\text{unscaled}}_{\alpha}(X_{n+1})\}\geq 1-\big(\alpha\mathbb{E}[\widehat{m}^{*}]+\varepsilon+\delta\big)\quad\forall\alpha\in(0,1).$
The prediction set from thresholding $\widehat{m}^{*}p_{\text{all}}$ is $\widehat{C}_{\alpha}^{\text{scaled}}(X_{n+1})=\{y:\widehat{m}^{*}p_{\text{all}}(X_{n+1},y)>\alpha\}$, and it satisfies the coverage guarantee
$\mathbb{P}\{Y_{n+1}\in\widehat{C}^{\text{scaled}}_{\alpha}(X_{n+1})\}\geq 1-(\alpha+\varepsilon+\delta)\quad\forall\alpha\in(0,1).$
Here, $\widehat{m}^{*}$ is the empirical correction factor defined in (9).
The two formulations in Propositions 5 and 5 demonstrate the complementary roles of $m^{*}$. In the unscaled sets, $m^{*}$ scales the coverage lower bound in parallel with Proposition 5, while in the scaled sets it rescales the p-variable to yield a standard $1-\alpha$ style guarantee. For consistency with other conformal methods, we use the scaled formulation in our experiments.
An important practical question is how the quality of the correction $\widehat{m}^{*}$ depends on the merging set $S_{\text{merge}}$. In §B.1.1, we study this dependence by varying $|S_{\text{merge}}|$, and we show that even a modest number of samples (<200<200) is sufficient for accurate coverage.
We summarize our full procedure in Algorithm 1.
@@
Data-dependent weights allow the influence of each model to be adjusted based on how well it performs for a specific data point. This approach enables the construction of aggregated prediction sets that are tailored to the characteristics of the given data, which can be viewed as a form of locally conditional coverage. To be clear, true X-conditional coverage—where the coverage guarantee is conditioned on the current input—is impossible without additional distributional assumptions [33, 34]. However, we demonstrate in §6 that data-dependent weights allow us to create aggregated prediction sets with data-adaptive coverage, which can greatly improve conditional validity in practice.
@@
(Note that (10) is not without precedent: for the $1-2\alpha$ coverage guarantees of aggregation methods like cross-conformal and jackknife+, $\alpha^{\prime}=0.5$ is the highest significance level of interest.)
@@ 6 Application: mixture-of-experts
Refer to captionRefer to captionRefer to captionRefer to caption
@@
The routing network’s ability to learn data-dependent weights for each expert makes MoE a natural setting for applying our method of aggregating prediction sets by weighted p-values. We collect the routing weights into the simplex-valued vector $W(x):=(W_{1}(x),\dots,W_{K}(x))$; then, for a new input $X_{n+1}$, the routing network outputs the weight vector $\mathbf{W}^{(n+1)}:=W(X_{n+1})$, whose $k$th component we denote $W^{(n+1)}_{k}$. Let $\widehat{p}_{k}$ denote the p-value function of the $k$th expert. The weighted aggregate p-value function for label $y$ is
@@
At significance level $\alpha$, we form the MoE prediction set by thresholding the corrected p-variable $\widehat{m}^{*}p_{\text{all}}$, as established in Proposition 5:
$\widehat{C}_{\alpha}^{\text{MoE}}(X_{n+1})=\{y\in\mathcal{Y}:\widehat{m}^{*}p_{\text{all}}(X_{n+1},y;\mathbf{W}^{(n+1)})>\alpha\}.$
Intuitively, the routing weights adapt the contribution of each expert’s p-value to the input so that experts with higher predictive relevance for $X_{n+1}$ have greater influence. This data-adaptive weighting yields coverage guarantees that are locally more robust than split conformal with the full MoE predictor (Figure 2).
@@
To evaluate local validity, we compare against conformal quantile regression (CQR) [35], a widely used locally adaptive method. A strength of our framework is that it complements, rather than competes with, other adaptive methods, allowing weighted aggregation to be layered on top of CQR. Accordingly, we also evaluate a hybrid method that combines weighted aggregation with CQR. We assess performance using marginal coverage, worst-slice (WS) coverage [36], and prediction set size.
@@
Main regression results
In Figure 3, we compare split conformal to weighted aggregation using both absolute residual scores and CQR scores. From the marginal coverage plots (left column), we see that split conformal consistently achieves coverage closest to the nominal $1-\alpha$ level, while the weighted aggregation methods tend to overcover. However, the WS coverage plots (middle column) expose a well-known limitation of split conformal: it does not maintain coverage across the data, systematically undercovering on the WS slab. CQR scores improve split conformal WS coverage on some datasets (Airfoil and Communities), but they are still insufficient to close the coverage gap.
Weighted aggregation significantly improves WS coverage, coming close to or meeting the nominal level on all datasets. In addition, the gap between marginal and WS coverage is notably smaller for weighted aggregation than for split conformal, indicating that weighted aggregation provides more uniform coverage, even over the challenging regions of the data. These results show that weighted aggregation is useful in applications where local validity is a priority.
@@ 7 Conclusion
The asymmetric (weighted) aggregation of prediction sets is a flexible generalization of standard symmetric aggregation. Intuitively, coverage in this setting should vary with the distribution of weights. The results of Vovk and Wang [18] allow us to formalize this intuition for conformal prediction, so that for data-independent weights based on expert priors, we obtain improved guarantees when the weights are sufficiently asymmetric. We extend this framework to data-dependent weights (e.g. weights learned from data), enabling adaptive coverage that reflects the observed input. Experiments on WS coverage and demographic subgroups confirm the practical benefits of this extension, showing that weighted aggregation yields more reliable coverage in challenging settings.
@@ Acknowledgements
We thank Josh Gleason and Uthsav Chitra for their thoughtful feedback, and Josh Gleason in particular for many insightful discussions throughout this work. GW and RC are partially supported by a ONR MURI grant N00014-20-1-2781. DP is partially supported by an Amazon AI2AI fellowship. DP, SS, and AL are partially supported by the Gordon and Betty Moore Foundation grant #12128. AL is also supported by an Amazon Research Award.
@@ A.1 Proof of Proposition 5
A.2 Proof of Proposition 5
We define $m^{*}$ to be the smallest positive value such that $m^{*}P_{\text{all}}$ is a p-variable, i.e.
$\mathbb{P}\left\{m^{*}P_{\text{all}}\leq\alpha\right\}\leq\alpha,$ (13)
or, equivalently,
$\mathbb{P}\left\{m^{*}P_{\text{all}}>\alpha\right\}\geq 1-\alpha.$
$\widehat{C}^{\text{scaled}}_{\alpha}(X_{n+1})=\left\{y\in\mathcal{Y}:m^{*}p_{\text{all}}(X_{n+1},y;\mathbf{W}^{(n+1)})>\alpha\right\}$
has coverage
$\mathbb{P}\left\{Y_{n+1}\in\widehat{C}^{\text{scaled}}_{\alpha}(X_{n+1})\right\}\geq 1-\alpha.$
Alternatively, we can rearrange (13) to be
@@ A.2 Proof of Proposition 5
$\mathbb{P}\left\{P_{\text{all}}\leq\alpha^{\prime}\right\}\leq m^{*}\alpha^{\prime}$
or
@@
Under exchangeability, the prediction set
$\widehat{C}^{\text{unscaled}}_{\alpha}(X_{n+1})=\left\{y\in\mathcal{Y}:p_{\text{all}}(X_{n+1},y;\mathbf{W}^{(n+1)})>\alpha\right\}$
satisfies
$\mathbb{P}\left\{Y_{n+1}\in\widehat{C}^{\text{unscaled}}_{\alpha}(X_{n+1})\right\}\geq 1-m^{*}\alpha.$
Note that other transformations of $P_{\text{all}}$ can also yield valid p-variables while preserving the relative proportion of weights. For example, if $P_{\text{all}}$ is shifted rather than scaled, then the condition
@@
also leads to coverage of at least $1-\alpha$ for the scaled set, and at least $1-(m^{\prime}+\alpha)$ for the unscaled set. We focus on the correction factor $m^{*}$ because it aligns with the framework of Vovk and Wang [18] and works well in practice.
A.3 Proof of Proposition 5
@@
$\mathbb{P}\left\{\widehat{m}^{*}P_{\text{all}}\leq\alpha\mid S_{\text{merge}}\right\}\leq\alpha+\varepsilon.$ (17)
Under exchangeability, the prediction set
$\widehat{C}^{\text{scaled}}_{\alpha}(X_{n+1})=\left\{y\in\mathcal{Y}:\widehat{m}^{*}p_{\text{all}}(X_{n+1},y;\mathbf{W}^{(n+1)})>\alpha\right\}$
@@
$\mathbb{P}\left\{Y_{n+1}\notin\widehat{C}^{\text{scaled}}_{\alpha}(X_{n+1})\mid S_{\text{merge}}\right\}\leq\alpha+\varepsilon.$
@@
$\text{MC}=\mathds{1}\left\{Y_{n+1}\notin\widehat{C}^{\text{scaled}}_{\alpha}(X_{n+1})\right\}.$
@@
$\leq\left(\alpha+\varepsilon\right)\mathbb{P}(\mathcal{G})+1\cdot\mathbb{P}(\mathcal{G}^{c})$
$\leq\alpha+\varepsilon+\delta.$
This gives the coverage guarantee
$\mathbb{P}\left\{Y_{n+1}\in\widehat{C}^{\text{scaled}}_{\alpha}(X_{n+1})\right\}\geq 1-\left(\alpha+\varepsilon+\delta\right).$
To derive the guarantee for the unscaled prediction set, we apply a change of variable to (17) to get
$\mathbb{P}\left\{P_{\text{all}}\leq\alpha^{\prime}\mid S_{\text{merge}}\right\}\leq\widehat{m}^{*}\alpha^{\prime}+\varepsilon.$
The remainder of the proof follows the same conditioning and marginalization argument, with the substitution of $\widehat{m}^{*}\alpha$ for $\alpha$ carried through. To be explicit, the prediction set
$\widehat{C}^{\text{unscaled}}_{\alpha}(X_{n+1})=\left\{y\in\mathcal{Y}:p_{\text{all}}(X_{n+1},y;\mathbf{W}^{(n+1)})>\alpha\right\}$
has a conditional miscoverage guarantee of
$\mathbb{P}\left\{Y_{n+1}\notin\widehat{C}^{\text{unscaled}}_{\alpha}(X_{n+1})\mid S_{\text{merge}}\right\}\leq\widehat{m}^{*}\alpha+\varepsilon.$
Then miscoverage indicator $\text{MC}^{\prime}=\mathds{1}\left\{Y_{n+1}\notin\widehat{C}^{\text{unscaled}}_{\alpha}(X_{n+1})\right\}$ can help us marginalize over $S_{\text{merge}}$:
$\mathbb{P}\left\{\text{MC}^{\prime}=1\right\}$ $=\mathbb{E}\left[\mathbb{P}\left\{\text{MC}^{\prime}=1\mid S_{\text{merge}}\right\}\right]$
@@
This gives the guarantee
$\mathbb{P}\left\{Y_{n+1}\in\widehat{C}^{\text{unscaled}}_{\alpha}(X_{n+1})\right\}\geq 1-\left(\alpha\mathbb{E}[\widehat{m}^{*}]+\varepsilon+\delta\right).$
Note that the same argument applies directly if we replace the empirical CDF $\widehat{F}_{P_{\text{all}}}$ by the conservative version $\widehat{F}_{P_{\text{all}}}^{\text{cons}}$ (8) because the two differ by at most $(|S_{\text{merge}}|+1)^{-1}$, a deterministic offset that can simply be absorbed into the DKW tolerance by replacing $\varepsilon$ with $\varepsilon+(|S_{\text{merge}}|+1)^{-1}$.
@@ A.5 Computing $\widehat{m}^{*}$ from an empirical CDF
$\lambda\in\{m>0:\mathbb{P}\{mP_{\text{all}}\leq\alpha\}\leq\alpha\penalty 10000\ \forall\alpha\in(0,1)\}=\{m>0:F_{P_{\text{all}}}\left(\alpha/m\right)\leq\alpha\penalty 10000\ \forall\alpha\in(0,1)\}.$
@@
$m^{*}=\inf\{m>0:F_{P_{\text{all}}}\left(\alpha/m\right)\leq\alpha\penalty 10000\ \forall\alpha\in(0,1)\}=\sup_{\delta>0}{\frac{F_{P_{\text{all}}}(\delta)}{\delta}}.$ (18)
@@
$\widehat{F}_{P_{\text{all}}}(\alpha)=\frac{\sum_{i\in S_{\text{merge}}}\mathds{1}\left\{p_{\text{all}}(X_{i},Y_{i};\mathbf{W}^{(i)})\leq\alpha\right\}}{|S_{\text{merge}}|},$
@@ B.1 Synthetic data
Figure 4 illustrates how the size of the merging set $|S_{\text{merge}}|$ impacts mean coverage. On the left plot, we use the $m^{\dagger}$ correction (10) for WA targeted (that is, $(0,\alpha)$ aggregation), and we compare different ways of assigning features to experts. On the right plot, we use a non-overlapping feature assignment, and we compare the different variants of weighted aggregation. Our results show that larger $|S_{\text{merge}}|$ leads to tighter coverage, with 160 samples being sufficient to achieve $3\%$ overcoverage for most feature assignment methods. Interestingly, configurations where experts have no overlapping features tend to overcover the most. We explore the effect of feature assignment in the next section.
@@ B.2.2 Local validity over demographic groups
Figures 8 and 9 compare split conformal with weighted aggregation variants in terms of coverage and prediction set size across demographic groups. Unlike Figures 3 and 7, where local validity is assessed via WS coverage, these experiments evaluate local validity in terms of consistency across demographic groups. Thus, we display group-specific performance for each method, which we group further by the two types of nonconformity score, absolute residual and CQR score.
The demographic groups in Figures 8 and 9 represent communities where a particular racial demographic is in the top $p$-percentile of representation [38]. Figure 8 shows results for $p=50$, and Figure 9 for $p=70$.